# Data Guard: A Fine-grained Purpose-based Access Control System for Large Data Warehouses

### v1

@@ 4.1. Data Access Policies
4.1.1. Purpose
The access control model described in this paper is motivated by the purpose limitation principle introduced in the GDPR regulation (GDPR, 2016). Purpose limitation requires that data is collected from data subjects (i.e. individuals whose personal data is collected) for an explicitly declared purpose and is only processed for use cases compatible with the purpose under which it is collected. Further, usage of an individual’s data for newer purposes is driven based on consents from the individual. At LinkedIn, a purpose is used to indicate the business justification for data collection or data processing. Purposes can either map to an external product such as advertising and jobs, or correspond to internal use cases such as security and business analytics.
At LinkedIn, purposes are organized as a directed acyclic graph which enables policy inheritance across purposes and minimizes policy duplication. For simplicity of discussion however, we assume henceforth that purposes are arbitrary string literals (such as ads and jobs) for the remainder of the paper.
@@ 4.1.1. Purpose
4.1.2. Policy Labels
Data producers provide semantic description of data in the tables they produce using policy labels. Figure 2 shows a data producer assigning policy labels $employer$ and $education$ assigned to $col_{1}$ and $col_{2}$ of table $T_{1}$. In general, data producers can assign policy labels either to: (i) a table, (ii) column in a table (either a top-level or a nested column), (iii) cells, or (iv) sub-cell-level data elements within arrays, structs and maps. As we will see in Section 4.3, this fine-grained assignment of policy labels allows Data Guard to apply data masking at much finer granularities than existing solutions.
Figure 2 also shows that policy labels are stored as metadata inside a Data Catalog. This design choice is motivated by recent trends in data architecture where data catalogs such as Horizon (Horizon, 2024), Unity (Unity, 2024) and DataHub (Datahub, 2019) have emerged as the foundational infrastructure component to address data discovery and data governance needs of large organizations. These catalogs serve as the inventory of the critical data assets (e.g. warehouse tables) of an organization and store a variety of metadata such as schemas, lineages and ownership associated with these assets. These catalogs are therefore well-suited for storing compliance metadata such as policy labels.
At LinkedIn, policy labels are carefully curated and organized into a hierarchical structure referred to as ontology. An ontology-based organization not only classifies data in our ecosystem, but also captures relationships between these data classes. Ontology also supports advanced use cases at LinkedIn related to reasoning and inference on compliance metadata. A detailed discussion on the design of our ontology is beyond the scope of this paper. We note however that the access control model proposed in this paper does not require an ontology-based organization of policy labels. For simplicity, we therefore assume that policy labels are arbitrary string values for the remainder of the paper.
@@ 4.1.3. Policy Language
Let $\mathcal{P}$ denote the set of purposes and $\mathcal{L}$ denote the set of policy labels. A policy in Data Guard is a tuple $(p,l,\langle c,a\rangle)$ where $p\in\mathcal{P}$ and $l\in\mathcal{L}$. $\langle c,a\rangle$ is a condition and action pair (or, a rule) where:
  • $c$ is a compound SQL predicate composed of simple predicates connected using the boolean operators $\{AND,OR,NOT\}$. Each simple predicate in $c$ is of the form $(x\circ y)$ where:
    1. (1)
      where $x\in\mathcal{X}$ is a set of attributes (defined below),
    2. (2)
      $y\in dom(x)$ is the domain of values of attribute $x$.
    3. (3)
      $\circ$ represents a relational operator from the set of relational operators $\{=,\neq,<,>,\leq,\geq,BETWEEN,IN,\\ LIKE,ISNULL\}$.
  • $a\in\mathcal{A}$ represents an action from the set of actions $\mathcal{A}$ that the access control system must take when $c$ is true.
Thus, the example policy shown in Figure 2 can be represented as a tuple: $(ads,employer,\langle allowEmplForAds=true,KEEP\rangle)$.
The set $\mathcal{X}$ of attributes over which policy conditions are defined is a union of data subject attributes (e.g. opt-in/opt-out consents, user location), data accessor attributes (e.g. accessor’s role and location), and system attributes (e.g. availability/security zone where the system is located in). The set of actions $\mathcal{A}$ of interest in this paper is the set $\{KEEP,MASK\}$, where the action $KEEP$ allows access to the target data while $MASK$ redacts the data. For simplicity, we only consider a specific form of data redaction in this paper, where the target data element is either replaced with $NULL$ value or the tuple containing the target data element is filtered from the result set. Assignment of non-null replacement values to target data elements is an area of ongoing work and will be discussed in greater detail in Section 7.
At LinkedIn, policies are managed like source code and policy changes are manually reviewed before they are committed to a code repository. Further, we run a number of validation checks on policies to avoid duplication and ensure that policies can be successfully parsed and compiled by the policy compiler. The changes committed to the repository are pushed to a Policy Catalog which provides APIs to query policies by purpose and policy labels.
Without loss of generality, all policies considered in the remainder of the paper are $KEEP$ policies i.e. the data element referenced by a policy label is preserved if the policy condition evaluates to true and masked otherwise. We further restrict our attention to policies based on data subject consents for the remainder of this paper. In doing so, we aim to highlight:
  1. (1)
@@
4.2. Data-Masking Views
Data Guard uses views as the interface for access control. Each data-masking view is a SQL representation of applicable data policies on a table for a given purpose. The applicable policies for a given purpose are found by matching policy labels assigned to fields in a table against the labels assigned to policies. The matched policies are then translated to SQL by the Policy Compiler, the implementation of which will be discussed in detail in Section 5.1.1.
We illustrate an example data-masking view created by Data Guard in Figure 2. Let us assume that the table $T_{1}$ shown in the figure is keyed by an $id$ column containing data subject identifiers. The figure also shows that $T_{1}$ contains a column $col_{2}$ that has $education$ data with an applicable policy for the $ads$ purpose. For simplicity, let us assume that all data subject attributes referenced in policy definitions are stored in the $member\_settings$ table which is also keyed by the data subject identifier. A naively implemented policy compiler generates the following view SQL for $ads$ purpose which masks $col_{2}$ based on the value of the attribute $allowGenderForAds$ in the $member\_settings$ table:
@@ 4.2. Data-Masking Views
Figure 3. A data-masking view for $ads$ purpose
The view SQL shown in Figure 3, while inefficient, has the important property that it is schema-preserving i.e. it has the same schema as the underlying table $T_{1}$. This invariant is maintained across all views created by Data Guard. The schema-preserving property ensures that applications can switch consumption from tables to views without needing to make code changes. Each view generated by Data Guard is registered within a purpose-specific database (e.g. $ads.V_{1}$ in Figure 2), which facilitates search and discovery of views in the data catalog. In Section 5, we discuss a number of optimizations which will be used to rewrite the view SQL in Figure 3 in order to make it performant.
@@
  • Agility: Implementing changes to policy compiler is much faster than making changes to the compute engine code, which caters to a much larger set of use cases beyond access control.
  • Optimizations: A dedicated policy compilation layer provides the flexibility of implementing custom optimizations (e.g. bitmap optimization discussed in Section 5.1.2) when generating view SQL. Such use-case specific optimizations are otherwise non-trivial to implement inside general-purpose compute engines.
@@
4.3. Field Paths
4.3.1. Motivation
Existing solutions (Stonebraker and Wong, 1974; Rizvi et al., 2004; Xue et al., 2023; Agrawal et al., 2005) support data masking at row, column and cell-level granularity. These masking operations are typically accomplished using traditional projection and selection operators. As mentioned in Section 1, Data Guard’s access control model supports masking data at much finer granularities than is possible with previous solutions. This need for finer grained data masking is motivated by prevalence of non-relational data in LinkedIn’s warehouse (and indeed, many modern data warehouses (Snowflake, 2024) and lakehouses (Databricks, 2024)). It is common to model data using relation-valued attributes and collection types such as arrays and maps. Masking data selectively inside such data structures cannot be done using relational operators such as projection and selection alone.
@@ 4.3.1. Motivation
$col_{2}$: STRUCT<$field_{21}$:BIGINT, $field_{22}$:VARCHAR>
$col_{3}$: ARRAY<STRUCT<$field_{31}$:VARCHAR, $field_{32}$:DOUBLE>>
$col_{4}$: MAP<VARCHAR, ARRAY<STRUCT<$field_{41}$:VARCHAR,
$field_{42}$:BOOLEAN$>$
Figure 4 shows an instance of $\mathcal{R}$.
@@
\$.$\$.col_{1}$
\$.[?(@.$col_{1}=^{\prime}def^{\prime})]$
\$.[?(@.$col_{1}=^{\prime}ghj^{\prime})].col_{2}$
\$.$\$.col_{3}.[item].[?(@.field_{31}=^{\prime}s1^{\prime})]$
\$.$\$.col_{4}.[value].[item].[?(@.field_{41}=^{\prime}v2^{\prime})].field_{42}$
Figure 4. Example of a nested relation. Data Guard’s field path expressions allow masking of data inside nested attributes such as structs, arrays and maps.
In order to mask data at a sub-cell granularity (e.g. $field_{31}$, $field_{42}$ shown in Figure 4), we need an operator for selecting this data. There have been prior proposals (Colby, 1990) which introduce recursive selection and projection operators, which can be used to select and mask data at sub-cell granularity. While there have been efforts to extend SQL language to support recursive operations (ISO-ANSI, 1999), very few commercial and open-source systems support such extensions. While introducing these extensions to existing open source systems like Spark and Trino is in theory possible, it would require a significant modification of the SQL standard adopted by these systems and a non-trivial investment of effort to drive alignment across the open-source community. On the other hand, there have been numerous path DSLs such as XMLPath (XPath 3.1, 2017) and JsonPath (Gössner et al., 2024) that have been developed and successfully adopted to support element-wise operators. Given these trade-offs, using a DSL to evaluate field paths efficiently during data access was a cost-effective alternative. Leveraging existing DSLs such as XPath and JsonPath was not an option due to several limitations that prevent their usage for structured data handling. XPath is intended exclusively for processing XML documents. JsonPath is schema-unaware and does not distinguish between types such as maps and structs. Nevertheless, we use JsonPath to guide the design of the field path DSL described next.
4.3.2. Field Path Expressions
@@
A field path is a sequence of operators $P_{1},P_{2},\ldots,P_{n}$, where each operator $P_{i}$ operates on the result set of $P_{i-1}$ and has one of the following types:
  1. (1)
    Root access operator: Denoted by a special symbol \$currency-dollar\\$\$, this operator is an identity operator whose result set is the input relation. The root access operator is always the first operator in the sequence of operators for any given field path.
  2. (2)
@@ 4.3.3. Examples
  1. (1)
    A dereference operator following a root access operator can be used to select an attribute from a relation. For example, \$.$\$.col_{1}$ selects $col_{1}$.
  2. (2)
    A filter operator following a root access operator is referred to as a row selector and has the ability to select rows from a relation. For example, the field path \$.[?(@.$col_{1}=^{\prime}def^{\prime})]$ selects rows from the relation satisfying $col_{1}=^{\prime}def^{\prime}$.
  3. (3)
    A row selector followed by a dereference operator selects cell values of an attribute from the selected rows. For example, the field path \$.[?(@.$col_{1}=^{\prime}ghj^{\prime})].col_{2}$ selects values of $col_{2}$ from rows satisfying the condition $col_{1}=^{\prime}ghj^{\prime}$.
  4. (4)
    A filter operator following an unnest operator on a collection type field (e.g. arrays and maps) filters elements of the collection. For example, the field path \$.$\$.col_{3}.[item].[?(@.field_{31}=^{\prime}s1^{\prime})]$ selects elements of $col_{3}$ such that each element satisfies the condition $field_{31}=^{\prime}s1^{\prime}$. Similarly, the field path \$.$\$.col_{4}.[value].[item].[?(@.field_{41}=^{\prime}v2^{\prime})].field_{42}$ selects values of the field $field_{42}$ from the collection of map values satisfying $field_{41}=^{\prime}v2^{\prime}$.
In summary, field path expressions allow data selection at much finer granularities than is possible using traditional selection and projection operators available in commercial databases.
@@
4.3.4. Data-Masking Operator
@@
(1) $mask(\mathcal{R},p,f)=\begin{cases}\sigma_{\neg\text{pred OR p.cond}}(\mathcal{R}),\text{if $f=\$.[?(pred)]$}\\ {\tau_{f\rightarrow m(f)}(\sigma_{\neg\text{p.cond}}(\mathcal{R})))}\cup\sigma_{\text{p.cond}}(\mathcal{R}),\text{otherwise}\end{cases}$
where $\tau$ denotes a transformation operator that masks the data element at path $f$ by applying a masking function $m$ and reassembles the transformed attribute (potentially, a nested relation) back into its original structure. Equation 1 shows that when $f$ does not contain a dereference operator, the masking operator reduces to a row-level mask and removes the matching rows from the result set. Otherwise, the transformation operator functions as a column-level mask applying the function $m$ to values of $f$. The masking function $m$ in Equation 1 has the following behavior:
@@
In summary, field paths and data masking operators give us the capability of masking data at a very fine granularity. The exact implementation of the data-masking operator will be discussed in detail in Section 5.1.1.
5. Implementation
5.1. View Creation
In this section, we provide a detailed description of how Data Guard translates access control policies into data-masking views. We also describe the various optimizations in the generated views that ensure data accesses through data-masking views are performant.
5.1.1. Policy Compiler
The process of translating data policies into SQL code is referred to as policy compilation. As shown in Algorithm 1, the policy compiler incrementally constructs a relational algebra plan for a given input relation $\mathcal{R}$ by iterating over a list of matching policies for $\mathcal{R}$ and then converting the plan into SQL text.
Algorithm 1 Compiling policies into SQL code where $enforce\_policy()$ function creates a relational algebra plan representing the enforcement of a data access policy $p$ on attribute $\mathcal{A}$ of the input relation $\mathcal{R}$, and $to\_sql()$ function returns the SQL text for a relational algebra plan
  Input: Input relation $\mathcal{R}$, a list $\mathcal{L}$ of matching pairs (field path, policy)
  Output: SQL string representing a data-masking view
  $plan\leftarrow\mathcal{R}$
  for $(f,p)\in\mathcal{L}$ do
     $plan\leftarrow enforce\_policy(plan,f,p)$
  end for
  return  $to\_sql(plan)$
A key step in Algorithm 1 is the call to $enforce\_policy()$ method to update the relational algebra plan. Under the hood, the $enforce\_\\ policy()$ method adds the data-masking operator $mask$ introduced in Section 4.3.4 to the plan. Thus, the efficiency of the generated plan (and of the resulting view SQL) depends crucially on the implementation of the $mask$ operator.

### v2

@@ IV-A Data Access Policies
IV-A1 Purpose
The access control model described in this paper is motivated by the purpose limitation principle introduced in the GDPR regulation [1]. Purpose limitation requires that data is collected from data subjects (i.e. individuals whose personal data is collected) for an explicitly declared purpose and is only processed for use cases compatible with the purpose under which it is collected. Further, usage of an individual’s data for newer purposes is driven based on consents from the individual. At LinkedIn, a purpose is used to indicate the business justification for data collection or data processing. Purposes can either map to an external product such as advertising and jobs, or correspond to internal use cases such as security and business analytics.
@@ IV-A1 Purpose
IV-A2 Policy Labels
Data producers provide semantic description of data in the tables they produce using policy labels. Figure 2 shows a data producer assigning policy labels employeremployer and educationeducation assigned to $col_{1}$ and $col_{2}$ of table $T_{1}$. In general, data producers can assign policy labels either to: (i) a table, (ii) column in a table (either a top-level or a nested column), (iii) cells, or (iv) sub-cell-level data elements within arrays, structs and maps. This fine-grained assignment of policy labels allows Data Guard to apply data masking at much finer granularities than existing solutions.
Figure 2 also shows that policy labels are stored as metadata inside a Data Catalog. This design choice is motivated by recent trends in data architecture where data catalogs such as Horizon [30], Unity [31] and DataHub [32] have emerged as the foundational infrastructure component to address data discovery and data governance needs of large organizations. These catalogs serve as the inventory of the critical data assets (e.g. warehouse tables) of an organization and store a variety of metadata such as schemas, lineages and ownership associated with these assets. These catalogs are therefore well-suited for storing compliance metadata such as policy labels.
@@ IV-A3 Policy Language
Let $\mathcal{G}$ denote the set of purposes and $\mathcal{L}$ denote the set of policy labels. A policy $p$ in Data Guard is a tuple $(g,l,\langle c,a\rangle)$ where $g\in\mathcal{G}$ and $l\in\mathcal{L}$. $\langle c,a\rangle$ is a condition and action pair (or, a rule) where:
  • $c$ is a compound SQL predicate composed of simple predicates connected using the boolean operators $\{\texttt{AND},\texttt{OR},\texttt{NOT}\}$. Each simple predicate in $c$ is of the form $(x\mathbin{\circ}y)$ where:
    1. 1.
      $x\in\mathcal{X}$ is a set of attributes (defined below),
    2. 2.
      $y\in\mathrm{dom}(x)$ is the domain of values of attribute $x$.
    3. 3.
      $\circ$ represents a relational operator from the set of relational operators $\{=,\neq,<,>,\leq,\geq,\texttt{BETWEEN},\texttt{IN},\texttt{LIKE},\texttt{IS NULL}\}$.
  • $a\in\mathbb{A}$ represents an action from the set of actions $\mathbb{A}$ that the access control system must take when $c$ is true.
Thus, the example policy shown in Figure 2 can be represented as a tuple:
$(ads,employer,\langle allowEmpForAds=true,\mathrm{KEEP}\rangle)$ (1)
The set $\mathcal{X}$ of attributes over which policy conditions are defined is a union of data subject attributes (e.g. opt-in/opt-out consents, user location), data accessor attributes (e.g. accessor’s role and location), and system attributes (e.g. availability/security zone where the system is located in). The set of actions $\mathbb{A}$ of interest in this paper is the set $\{\texttt{KEEP},\texttt{MASK}\}$, where the action KEEP allows access to the target data while MASK redacts the data. For simplicity, we only consider a specific form of data redaction in this paper, where the target data element is either replaced with NULL value or the tuple containing the target data element is filtered from the result set. Note that the same model can be extended to define and support other types of actions like truncating, or partially masking the data.
Without loss of generality, all policies considered in the remainder of the paper are KEEP policies i.e. the data element referenced by a policy label is preserved if the policy condition evaluates to true and masked otherwise. We further restrict the policy conditions to be based on data subject consents only. This restriction is motivated by the fact that data subject consents are the most common basis for data access policies at LinkedIn. We therefore focus our attention to those consent-based policies for the remainder of this paper.
In doing so, we aim to highlight:
  1. 1.
@@
IV-B Data-Masking Views
Data Guard uses views as the interface for access control. Each data-masking view is a SQL representation of applicable data policies on a table for a given purpose. The applicable policies for a given purpose are found by matching policy labels assigned to fields in a table against the labels assigned to policies. The matched policies are then translated to SQL by the Policy Compiler, the implementation of which will be discussed in detail in Section V-A1.
We illustrate an example data-masking view created by Data Guard in Figure 2. Let us assume that the table $T_{1}$ shown in the figure is keyed by an $id$ column containing data subject identifiers. The figure also shows that $T_{1}$ contains a column $col_{2}$ that has educationeducation data with an applicable policy for the adsads purpose. For simplicity, let us assume that all data subject attributes referenced in policy definitions are stored in the member_settingsmember\_settings table which is also keyed by the data subject identifier. A naively implemented policy compiler generates the following view SQL for adsads purpose which masks $col_{2}$ based on the value of the attribute $allowGenderForAds$ in the member_settingsmember\_settings table:
@@ IV-B Data-Masking Views
Figure 3: A data-masking view for adsads purpose
The view SQL shown in Figure 3, while inefficient, has the important property that it is schema-preserving i.e. it has the same schema as the underlying table $T_{1}$. This invariant is maintained across all views created by Data Guard. The schema-preserving property ensures that applications can switch consumption from tables to views without needing to make code changes. Each view generated by Data Guard is registered within a purpose-specific database (e.g. $ads.V_{1}$ in Figure 2), which facilitates search and discovery of views in the data catalog. In Section V, we discuss a number of optimizations which will be used to rewrite the view SQL in Figure 3 in order to make it performant.
@@
  • Agility: Because the policy compiler is independent of the compute engine, implementing changes there—such as generating optimized data-masking views—is faster and less risky than modifying the compute engine, which must support many other use cases beyond access control.
  • Optimizations: A dedicated policy compilation layer provides the flexibility of implementing custom optimizations (e.g. bitmap optimization discussed in Section V-A2) when generating view SQL. Such use-case specific optimizations are otherwise non-trivial to implement inside general-purpose compute engines.
@@
IV-C Field Paths
IV-C1 Motivation
Existing solutions [12, 13, 8, 14] support data masking at row, column and cell-level granularity. These masking operations are typically accomplished using traditional projection and selection operators. As mentioned in Section I, Data Guard’s access control model supports masking data at much finer granularities than is possible with previous solutions. This need for finer grained data masking is motivated by prevalence of non-relational data in LinkedIn’s warehouse (and indeed, many modern data warehouses [19] and lakehouses [20]). It is common to model data using relation-valued attributes and collection types such as arrays and maps.
In many ML applications, user attributes and features are stored in a structured yet sparse format, such as arrays of key-value pairs or nested feature maps, avoiding the need for wide and constantly evolving tables as new features are added. Additionally, some tables combine data from different sources in different rows, or as different elements in arrays. The ability to represent, label and mask data within arrays or maps conditionally provides fine-grained control over compliance while preserving the efficiency and convenience of representation. Masking data selectively inside such data structures cannot be done using relational operators such as projection and selection alone.
@@ IV-C1 Motivation
[…]
$field_{42}$:BOOLEAN>>>
@@
$\mathdollar.col_{1}$
$\mathdollar.[?(@.col_{1}=^{\prime}def^{\prime})]$
$\mathdollar.[?(@.col_{1}=^{\prime}ghj^{\prime})].col_{2}$
$\mathdollar.col_{3}.[item].[?(@.field_{31}=^{\prime}s1^{\prime})]$
Figure 4: Example of a nested relation with arrays, structs, maps.
In order to mask data at a sub-cell granularity (e.g. $field_{31}$, $field_{42}$ shown in Figure 4), we need an operator for selecting this data. There have been prior proposals [33] which introduce recursive selection and projection operators, which can be used to select and mask data at sub-cell granularity. While there have been efforts to extend SQL language to support recursive operations [34], very few commercial and open-source systems support such extensions. While introducing these extensions to existing open source systems like Spark and Trino is in theory possible, it would require a significant modification of the SQL standard adopted by these systems and a non-trivial investment of effort to drive alignment across the open-source community. On the other hand, there have been numerous path DSLs such as XMLPath [35] and JsonPath [36] that have been developed and successfully adopted to support element-wise operators. Given these trade-offs, using a DSL to evaluate field paths efficiently during data access was a cost-effective alternative. Leveraging existing DSLs such as XPath and JsonPath was not an option due to several limitations that prevent their usage for structured data handling. XPath is intended exclusively for processing XML documents. JsonPath is schema-unaware and does not distinguish between types such as maps and structs. Nevertheless, we use JsonPath to guide the design of the field path DSL described next.
IV-C2 Field Path Expressions
@@
A field path is a sequence of operators $P_{1}$, $P_{2}$, …, $P_{n}$, where each operator $P_{i}$ operates on the result set of $P_{i-1}$ and has one of the following types:
  1. 1.
[…]
  2. 2.
@@ IV-C3 Examples
  1. 1.
    A dereference operator following a root access operator can be used to select an attribute from a relation. For example, $\mathdollar.col_{1}$ selects $col_{1}$.
  2. 2.
    A filter operator following a root access operator is referred to as a row selector and has the ability to select rows from a relation. For example, the field path $\mathdollar.[?(@.col_{1}=\text{`def'})]$ selects rows from the relation satisfying $col_{1}=\text{`def'}$.
  3. 3.
    A row selector followed by a dereference operator selects cell values of an attribute from the selected rows. For example, the field path $\mathdollar.[?(@.col_{1}=\text{`ghj'})].col_{2}$ selects values of $col_{2}$ from rows satisfying the condition $col_{1}=\text{`ghj'}$.
@@
IV-C4 Data-Masking Operator
@@
$\small mask(\mathcal{R},p,f)=\begin{cases}\sigma_{\neg\text{pred OR p.c}}(\mathcal{R}),\qquad\text{if }f=\mathdollar.[?(pred)]\\ {\tau_{f\rightarrow m(f)}(\sigma_{\neg\text{p.c}}(\mathcal{R})))}\cup\sigma_{\text{p.c}}(\mathcal{R}),\text{otherwise}\end{cases}$ (2)
where $\sigma$ is the standard relational algebra select operator. $\tau$ denotes a transformation operator that masks the data element at path $f$ by applying a masking function $m$ and reassembles the transformed attribute (potentially, a nested relation) back into its original structure. Equation 2 shows that when $f$ does not contain a dereference operator, the masking operator reduces to a row-level mask and removes the matching rows from the result set. Otherwise, the transformation operator functions as a column-level mask applying the function $m$ to values of $f$. The masking function $m$ in Equation 2 has the following behavior:
@@
In summary, field paths and data masking operators give us the capability of masking data at a very fine granularity. The exact implementation of the data-masking operator will be discussed in detail in Section V-A1.
V Implementation
V-A View Creation
In this section we first present the Policy Compiler (Section V-A1), which translates access-control policies into data-masking views. This section describes several optimizations (Sections V-A2 and V-A3) that ensured scalability of data-masking views in production. While these optimizations are presented as heuristics in this paper, we plan to formalize them into a general optimization framework in future work. We conclude by describing the view maintenance system that handles policy and label changes (Section V-A4).
V-A1 Policy Compiler
The process of translating data policies into SQL code is referred to as policy compilation. A policy compiler takes as input a relation $\mathcal{R}$ (table or view) and a list $\mathcal{L}$ of matching pairs (field path, policy) for $\mathcal{R}$. The output of the policy compiler is a SQL string representing a data-masking view for $\mathcal{R}$.